# ImpMIA: Leveraging Implicit Bias for Membership Inference Attack under Realistic Scenarios

### v1

@@ 4 ImpMIA Attack
In this section, we introduce ImpMIA, our white-box membership inference attack that exploits the implicit bias of neural networks. The attack builds on the observation from Eq. 3 that trained parameters can be represented as a linear combination of per-sample gradients from the training samples (members). Thus, members can be distinguished from non-members by their relative contribution to this representation of the parameters, where members contribute to this reconstruction while non-members do not. We first outline the practical construction of the attack in Section 4.1, and then detail the technical optimization procedures and stabilization strategies in Section 4.2
@@ 4.1 Practical Attack Construction
In practice, the attacker does not know the true training set, but only has access to a candidate pool $X_{\mathrm{sup}}=\{(x_{i},y_{i})\}_{i=1}^{M}$, which contains an unknown subset of training samples $X_{\mathrm{train}}$ and non-members $X_{\mathrm{test}}$. However, we can still optimize the coefficients using all samples, deriving a $\lambda$ coefficient for each (either member or not). This provides the key signal: we expect the coefficients of training samples to be significantly larger, while those of non-members remain small. This is because the number of network parameters is typically much larger than the size of the candidate pool, and therefore deriving the correct weights in the optimization is much more likely when true members exert stronger influence. Importantly, when $|X_{\mathrm{sup}}|$ is smaller than the dimension of $\theta$ and the vectors $\{g_{i}\}$ are linearly independent, the system admits a unique solution for $\{\lambda_{i}\}$. Note that, following Eq. 2&3, we expect large coefficients for training samples near the margin (close to the decision boundary), while other samples are expected to have lower coefficients (zero in theory).
@@
and stack these into the matrix $A=[g_{1}\ |\ \cdots\ |\ g_{M}]\in\mathbb{R}^{p\times M},$
where $p$ is the number of model parameters. If the training set were known, we could restrict to $A_{\mathrm{train}}$ and solve exactly for multipliers $\lambda_{\mathrm{train}}$ such that $A_{\mathrm{train}}\,\lambda_{\mathrm{train}}=\theta$. Since the training set is unknown, we instead solve the full system $A\,\lambda_{\text{score}}=\theta,$ deriving $M$ (number of candidates) coefficients. The resulting coefficients are used to calculate a membership score, where large values for a specific sample are interpreted as evidence of a higher probability that this sample was part of the training data (i.e., a member). To improve robustness and suppress spurious large values from non-members, we incorporate additional techniques for regularization and aggregation (detailed in Section 4.2). In Figure 3, we present a scatter plot of members and non-members, where the y-axis shows the $\lambda$ score and the x-axis the sample’s distance from the margin (i.e., $\Phi_{y_{i}}(x_{i};\theta)-\max_{j\neq y_{i}}\Phi_{j}(x_{i};\theta)$). As illustrated, high $\lambda$ scores are strong indicators of membership.
@@ 4.2 Implementation Details
Given a set of training samples and their corresponding classes, we optimize the $\lambda$ coefficients to satisfy the KKT conditions. We first filter out misclassified samples, since training members are more likely to be correctly predicted. For each remaining sample, we include both the original and its horizontal flip, reflecting the fact that most models are trained with augmentations, while keeping computational cost low by not adding further augmentations.
@@
Figure 2: Lambda Scores Visualization. Scatter plot of superset samples, with the x-axis showing distance from the decision boundary and the y-axis showing $\lambda$ scores; points are colored by membership (member vs. non-member). High $\lambda$ strongly indicates membership.
Figure 3: Effect of Assumption Removal. Performance of the best prior method (LiRA) and our method (ImpMIA) under the progressive removal of assumptions on CINIC-10. LiRA degrades sharply as assumptions are removed, while ImpMIA maintains stable performance.
@@ 5 Experimental Setting
We evaluate ImpMIA against both black-box and white-box baselines across CIFAR-10, CIFAR-100, and CINIC-10, using a standard ResNet-18 as the target model. Our experiments are designed to reflect realistic adversarial conditions, where the attacker is given only the trained model weights and a superset of candidate samples, with no knowledge of the training hyperparameters, data distribution, or member ratio. We first describe the datasets, evaluation metrics, and the competing baseline models (Section 5.1). We then detail the different scenarios corresponding to the assumptions usually made by reference-model–based MIAs, beginning with the setting where all assumptions are provided, and then moving to a more realistic setting without these assumptions (Section 5.2).
@@ 5.1 Models, Datasets and Evaluation Metrics
For all experiments, the target model is a ResNet-18 trained following the standard recipe of Cohen & Giryes (2024). Specifically, we use a batch size of $100$, a learning rate of $0.1$, momentum $0.9$ with Nesterov acceleration, weight decay of $10^{-4}$, and train for $400$ epochs using stochastic gradient descent (SGD) with standard data augmentations (random crop and horizontal flip). For each dataset and scenario, we randomly sampled 5 different sets of training samples, and trained 5 target models. Results are averaged across them.
We report performance using both aggregate metrics (AUC) and stringent low–false-positive criterion, True Positive Rate (TPR) at False Positive Rate (FPR) of 0.01% and 0.00%. This criterion introduced in Carlini et al. (2022), which observed that average-case metrics such as accuracy or AUC can be misleading: an attack may appear strong overall, yet fail completely in the regime of low false positives, which is the regime most relevant for privacy auditing (see Appendix E for a detailed discussion). In practice, an adversary cannot afford a large number of false alarms, as even a tiny FPR may translate to thousands of incorrectly flagged samples in real-world deployments. Therefore, evaluating TPR at very low FPR provides a stricter and more meaningful measure of membership inference risk.
We compare ImpMIA against recent state-of-the-art black-box and white-box membership inference attacks. For black-box attacks, we include Attack-P (Li & Zhang, 2021), Attack-R (Ye et al., 2022), and focused on the online version of LiRA (Carlini et al., 2022) and RMIA (Zarifzadeh et al., 2023), which currently represent the strongest black-box MIAs. For black-box baselines, we followed their best setting (e.g. training 256 reference models per experiment). For white-box attacks, we evaluate the adaptive self-influence attack (AdaSIF) (Cohen & Giryes, 2024), which represents the current state of the art in this category (see Section C.2 for additional implementation details). We also include as a baseline a simple white-box attack based on the magnitude of the network gradients, since Nasr et al. (2019) noted that this quantity provides the main signal in their attack (see Section C.2 for additional details). We were unable to compare with Nasr et al. (2019) and Leino & Fredrikson (2020), as their code is not available.
@@ 5.2 Membership Inference Attack Scenarios
The basic scenario is the one commonly assumed in black-box attacks, which remain the strongest-performing methods even when compared to existing white-box approaches. In the standard black-box setting, the attacker is assumed to know both the model architecture and the training hyperparameters used to train the target model (e.g., learning rate, optimizer, number of epochs). The attacker is also given a pool of samples containing both members (training samples) and non-members, under the assumptions that (i) the pool is drawn from the same distribution as the training data and (ii) the member/non-member ratio is known, often fixed at 1:1.
@@
These assumptions give reference-model attacks such as LiRA and RMIA a strong advantage: by training many reference models under the same conditions, they can effectively reconstruct the member/non-member loss distributions of the target model. This explains their strong performance in this regime, though it comes at the cost of training large model ensembles and relies on knowledge rarely available in practice.
@@
To evaluate attack performance under more realistic conditions, we carefully design experimental scenarios that relax these assumptions. In Section 6, we present results in the more realistic setting where those assumptions are not made, as well as in each individual scenario where one assumption is removed. Importantly, to ensure fair comparison across scenarios and with prior work, we kept the target model fixed and followed the basic scenario used in previous studies.
The assumption-elimination scenarios are:
  • Unknown Training Configuration – Since the attacker is not exposed to the target model’s training parameters, we trained the reference models of methods that require them using different settings. Specifically, for the reference models we used a different batch size (200 instead of 100), learning rate (0.01 instead of 0.1), weight decay ($10^{-3}$ instead of $10^{-4}$), and epochs (100 instead of 400) than those used for the ResNet-18 target model (detailed in Section 5.1).
  • Different Data Distribution – In this case, the attacker’s candidate pool mixes in-distribution data (the distribution from which the target model was trained) and out-of-distribution (OOD) data. For each dataset, we construct a pool of $50k$ samples by combining $30k$ in-distribution images with $20k$ OOD images (taken from another dataset). The target model is trained on $25k$ in-distribution samples drawn from the $30k$ portion, allowing for 5 repetitions of the experiment (with 5 different target models). In the reference-based attacks under the relevant online setting, the attacker trains each reference model on half of the superset, i.e., $25k$ examples sampled from the full mixed pool. OOD sources include a subset of ImageNet adapted to CIFAR-10 (CINIC-10 (Darlow et al., 2018)) and an enriched OpenImages dataset (Kuznetsova et al., 2020).
  • Removing All assumptions – Combining the above cases, the candidate pool for each dataset has $80k$ samples formed by mixing $30k$ in-distribution with $50k$ OOD images. The attacker trains reference models on a $40k$-example subset sampled from the full mixed pool under different configurations and without access to the true member ratio.
6 Results
In this section, we quantitatively compare our ImpMIA attack against prominent prior black-box and white-box attacks across multiple datasets. In Section 6.1, we demonstrate the superiority of our attack in the realistic scenario where only the model weights are known and the attacker is given a superset of samples that includes the training data, but without any additional knowledge of the model training or the superset composition. Next, in Section 6.2, we systematically analyze the impact of eliminating each assumption across different models, showing that reference-model–based methods suffer significant performance drops, while our method remains unaffected. Further analysis of the method and ablation studies can be found in Appendix B.
6.1 Membership Inference Results
We report membership inference attack results across three datasets, demonstrating the superiority of our proposed attack in the realistic scenario described above (without assumptions on model training, data distribution, or member ratio). As shown in Table 1, reference-model attacks such as LiRA and RMIA struggle severely in this regime: at 0% False Positive Rate (FPR), LiRA achieves only 0.17% True Positive Rate (TPR) on CIFAR-10, and RMIA 0.01% . Importantly, these are the best-performing attacks—even compared to white-box methods—when the knowledge assumptions are provided (see Table T4 in the appendix). Their effectiveness relies on training large sets of reference models under matching conditions, a requirement that breaks down when the attacker cannot replicate the target’s training process. In contrast, ImpMIA avoids training reference models entirely, and maintains strong TPR at low FPR across all datasets. On CIFAR-10, our attack achieves 1.41% TPR at 0% FPR, and 2.76% at 0.01% FPR, with similar or larger gains on the other datasets. Our approach achieves substantially stronger performance at strict low-FPR operating points, which are the most relevant in practical membership inference scenarios, and comparable AUC results. We further present the FPR–TPR trade-off curves (ROC curves) for all three datasets in Figure S2.
Table 1: Membership Inference Results. Performance of ImpMIA compared to both black-box (LiRA, RMIA) and white-box baselines across three datasets under the realistic no-assumptions setting. The main relevant metrics are TPR values at fixed false-positive rates (FPR = 0.00% and 0.01%), which capture detection power under stringent error constraints. ImpMIA significantly surpasses all other methods by a wide margin, due to their reliance on the different assumptions. For completeness, we also report AUC as an aggregate measure.
@@ 6.1 Membership Inference Results
Attack AUC @0.01 % @0.00 % AUC @0.01 % @0.00 % AUC @0.01 % @0.00 %
Attack-P 0.76 0.02 0.00 0.89 0.01 0.00 0.83 0.01 0.00
Attack-R 0.74 0.23 0.04 0.95 0.52 0.04 0.83 0.31 0.00
@@
GradNorm – margin 0.72 0.02 0.00 0.81 0.02 0.01 0.77 0.03 0.01
AdaSIF 0.80 0.05 0.00 0.92 0.01 0.00 0.85 0.01 0.00
@@
Table 2: Evaluation of Assumptions Influence. CINIC-10 membership inference across five scenarios: standard setting (all assumptions), removal of each assumption individually, and removal of all three simultaneously. Each entry shows TPR (%) at 0.01% / 0.00% FPR. The last column reports relative % of performance drop with respect to the Standard Setting.
Method
Standard
Setting
Unknown
Training Config.
Different
Distribution
Unknown Fraction
of Members
No Assumptions
TPR @ 0.01 / 0.00%
Attack-R 4.62 / 1.81 1.62 / 0.37 3.27 / 1.42 2.42 / 0.00 0.31 -93.3% / 0.00 -100%

### v3

@@ 4 ImpMIA Attack
In this section, we introduce ImpMIA, our white-box membership inference attack that exploits the implicit bias of neural networks. The attack builds on the observation from Eq. 5 that trained parameters can be represented as a linear combination of per-sample gradients from the training samples (members). Thus, members can be distinguished from non-members by their relative contribution to this representation of the parameters, where members contribute to this reconstruction while non-members do not. We first outline the practical construction of the attack in  Sec. 4.1, and then detail the technical optimization procedures and stabilization strategies in Sec. 4.2.
@@ 4.1 Practical Attack Construction
In practice, the attacker does not know the true training set, but only has access to a candidate pool $X_{\mathrm{sup}}=\{(x_{i},y_{i})\}_{i=1}^{M}$, which contains an unknown subset of training samples $X_{\mathrm{train}}$ and non-members $X_{\mathrm{test}}$. However, we can still optimize the coefficients using all samples, deriving a $\lambda$ coefficient for each (either member or not). This provides the key signal: we expect the coefficients of training samples to be significantly larger, while those of non-members remain small. This is because the number of network parameters is typically much larger than the size of the candidate pool, and therefore deriving the correct weights in the optimization is much more likely when true members exert stronger influence. Importantly, when $|X_{\mathrm{sup}}|$ is smaller than the dimension of $\theta$ and the vectors $\{g_{i}\}$ are linearly independent, the system admits a unique solution for $\{\lambda_{i}\}$. Note that, following Eqs. (4)-(5), we expect large coefficients for training samples near the margin (close to the decision boundary), while other samples are expected to have lower coefficients (zero in theory).
Figure 2: Lambda scores visualization. Scatter plot of evaluated samples, with the x-axis showing distance to the decision boundary and the y-axis showing $\lambda$ scores; points are colored by membership (member vs. non-member). High $\lambda$ indicates membership.
@@
and stack these into the matrix $A=[g_{1}\ |\ \cdots\ |\ g_{M}]\in\mathbb{R}^{p\times M},$ where $p$ is the number of model parameters. If the training set were known, we could restrict to $A_{\mathrm{train}}$ and solve exactly for multipliers $\lambda_{\mathrm{train}}$ such that $A_{\mathrm{train}}\,\lambda_{\mathrm{train}}=\theta$. Since the training set is unknown, we instead solve the full system $A\,\lambda_{\text{score}}=\theta,$ deriving $M$ (number of candidates) coefficients. The resulting coefficients are used to calculate a membership score, where large values for a specific sample are interpreted as evidence of a higher probability that this sample was part of the training data (i.e., a member). To improve robustness and suppress spurious large values from non-members, we incorporate additional techniques for regularization and aggregation (detailed in Section 4.2). In Fig. 2, we present a scatter plot of members and non-members, where the y-axis shows the $\lambda$ score and the x-axis the sample’s distance from the margin (i.e., $\Phi_{y_{i}}(x_{i};\theta)-\max_{j\neq y_{i}}\Phi_{j}(x_{i};\theta)$). As illustrated, high $\lambda$ scores are strong indicators of membership.
@@ 4.2 Implementation Details
Given an evaluation set of labeled samples, we optimize the $\lambda$ coefficients to satisfy the KKT conditions. We first filter out misclassified samples, since training members are more likely to be correctly predicted. For each remaining sample, we include both the original and its horizontal flip, reflecting the fact that most models are trained with augmentations, while keeping computational cost low by not adding further augmentations.
@@
Figure 3: TPR-FPR in the No-Auxiliary-Knowledge setting (CIFAR-10). We evaluate a setting with (i) unknown training configuration, (ii) distribution shift in the candidate pool, and (iii) unknown member ratio.
@@ 5 Experimental Setting
We evaluate ImpMIA against both black-box and white-box baselines across CIFAR-10, CIFAR-100, and CINIC-10, using a standard ResNet-18 as the target model. Our experiments reflect realistic adversarial conditions, where the attacker is given only the trained model weights and a superset of the training data, with no knowledge of the training hyperparameters, data distribution, or member ratio. We first describe the datasets, evaluation metrics, and the competing baseline models (Sec. 5.1). We then detail the different scenarios corresponding to the assumptions usually made by reference-model-based MIAs, beginning with the setting where all assumptions are provided, and then moving to a more realistic setting without them (Sec. 5.2).
@@ 5.1 Models, Datasets and Evaluation Metrics
For all experiments, the target model is a ResNet-18 trained following the standard recipe of Cohen and Giryes (2024). Specifically, we use a batch size of $100$, a learning rate of $0.1$, momentum $0.9$ with Nesterov acceleration, weight decay of $10^{-4}$, and train for $400$ epochs using stochastic gradient descent (SGD) with standard data augmentations (random crop and horizontal flip). For each dataset and scenario, we randomly sampled 5 different sets of training samples, and trained 5 target models. Results are averaged across them.
We report performance using both aggregate metrics (AUC) and stringent low-false-positive criterion, True Positive Rate (TPR) at False Positive Rate (FPR) of 0.01% and 0.0%. This criterion was introduced in Carlini et al. (2022), which observed that average-case metrics such as accuracy or AUC can be misleading: an attack may appear strong overall, yet fail completely in the regime of low false positives, which is the regime most relevant for privacy auditing (see Appendix E for a detailed discussion). In practice, an adversary cannot afford a large number of false alarms, as even a tiny FPR may translate to thousands of incorrectly flagged samples in real-world deployments. Therefore, evaluating TPR at very low FPR provides a stricter and more meaningful measure of membership inference risk.
We compare ImpMIA against recent state-of-the-art black-box and white-box membership inference attacks. For black-box attacks, we include Attack-P (Li and Zhang, 2021), Attack-R (Ye et al., 2022), and focused on the online version of LiRA (Carlini et al., 2022) and RMIA (Zarifzadeh et al., 2023), which currently represent the strongest black-box MIAs. For black-box baselines, we followed their best setting (e.g. training 256 reference models per experiment). For white-box attacks, we evaluate the adaptive self-influence attack (AdaSIF) (Cohen and Giryes, 2024), which represents the current state of the art in this category (see Sec. C.2 for additional implementation details). We also include as a baseline a simple white-box attack based on the magnitude of the network gradients, since Nasr et al. (2019) noted that this quantity provides the main signal in their attack (see Sec. C.2 for additional details). We were unable to compare with Nasr et al. (2019) and Leino and Fredrikson (2020), as their code is not available.
@@ 5.2 Membership Inference Attack Scenarios
The basic scenario is the one commonly assumed in black-box attacks, which remain the strongest methods even relative to existing white-box approaches. In the Full Auxiliary-Knowledge setting, the attacker is assumed to know both the model architecture and the training hyperparameters used to train the target model (e.g., learning rate, optimizer, epochs). The attacker is also given a pool of samples containing both members and non-members, under the assumptions that (i) the pool is drawn from the same distribution as the training data and (ii) the member/non-member ratio is known, often fixed at 1:1.
@@
These assumptions give reference-model attacks (e.g., LiRA, RMIA) a strong advantage: by training many reference models under identical conditions, they can reconstruct the target model’s member/non-member loss distributions. This explains their strong performance in this regime, though it comes at the cost of training large model ensembles and relies on knowledge rarely available in practice.
@@
To evaluate attack performance under more realistic conditions, we design experimental scenarios that relax these assumptions. In Sec. 6, we present results in the more realistic setting where those assumptions are not made, as well as in each individual scenario where one assumption is removed. Importantly, to ensure fair comparison across scenarios and with prior work, we kept the target model fixed and followed the basic scenario used in previous studies.
@@
Attack AUC @0.01 % @0.0 % AUC @0.01 % @0.0 % AUC @0.01 % @0.0 %
Attack-P 0.76 0.02 0.0 0.89 0.01 0.0 0.83 0.01 0.0
Attack-R 0.74 0.23 0.04 0.95 0.52 0.04 0.83 0.31 0.0
@@
GradNorm – margin 0.72 0.02 0.0 0.81 0.02 0.01 0.77 0.03 0.01
AdaSIF 0.80 0.05 0.0 0.92 0.01 0.0 0.85 0.01 0.0
@@
Table 1: Membership Inference results. Performance of ImpMIA compared to black-box (LiRA, RMIA) and white-box baselines across 3 datasets under the No-Auxiliary-Knowledge setting. The main relevant metrics are TPR values at fixed false-positive rates (FPR = 0.0% and 0.01%), which capture detection power under stringent error constraints. ImpMIA significantly surpasses all other methods by a wide margin, due to their reliance on the different assumptions. For completeness, we also report AUC as an aggregate measure.
The assumption-elimination scenarios are:
  • Unknown Training Configuration – Since the attacker is not exposed to the target model’s training parameters, we trained the reference models of methods that require them using different settings. Specifically, we used a different batch size (200 instead of 100), learning rate (0.01 instead of 0.1), weight decay ($10^{-3}$ instead of $10^{-4}$), and epochs (100 instead of 400) than those used for the ResNet-18 target model (detailed in Sec. 5.1).
  • Different Data Distribution – In this case, the attacker’s candidate pool mixes in-distribution data (the distribution from which the target model was trained) and out-of-distribution (OOD) data. For each dataset, we construct a pool of $50k$ samples by combining $30k$ in-distribution images with $20k$ OOD images (taken from another dataset). The target model is trained on $25k$ in-distribution samples drawn from the $30k$ portion, allowing for 5 repetitions of the experiment (with 5 different target models). In the reference-based attacks under the relevant online setting, the attacker trains each reference model on half of the evaluation set, i.e., $25k$ examples sampled from the full mixed pool. OOD sources include a subset of ImageNet adapted to CIFAR-10 (CINIC-10 (Darlow et al., 2018)) and an enriched OpenImages dataset (Kuznetsova et al., 2020).
  • No-Auxiliary-Knowledge Setting – Combining the above cases, the candidate pool for each dataset has $80k$ samples formed by mixing $30k$ in-distribution with $50k$ OOD images. The attacker trains reference models on a $40k$-example subset sampled from the full mixed pool under different configurations and without access to the true member ratio.
VGG16 ResNet50
Attack AUC @0.01 % @0.0 % AUC @0.01 % @0.0 %
LiRA 0.75 0.62 0.18 0.79 1.02 0.26
RMIA 0.78 0.69 0.06 0.78 0.09 0.02
Ours 0.78 1.59 0.56 0.82 1.80 0.78
Table 2: Membership Inference results on different architectures (No-Auxiliary-Knowledge setting) . Performance of LiRA, RMIA, and our method on VGG16 and ResNet50. We report AUC and TPR at fixed false-positive rates (FPR = 0.01% and 0.0%).
6 Results
In this section, we quantitatively compare our ImpMIA attack against prominent prior black-box and white-box attacks across multiple datasets. In Sec. 6.1, we demonstrate the superiority of our attack in the realistic scenario where only the model weights are known and the attacker is given a superset of samples that includes the training data, but without any additional knowledge of the model training or the superset composition. Next, in Sec. 6.2, we systematically analyze the impact of eliminating each assumption across different models, showing that reference-model-based methods suffer significant performance drops, while our method remains unaffected.
Method Full Auxiliary-Knowledge Unknown Training Config. Different Distribution Unknown Fraction of Members NO Auxiliary-Knowledge
TPR @ 0.01 / 0.0% Attack-R 4.62 / 1.81 1.62 / 0.37 3.27 / 1.42 2.42 / 0.0 0.31 -93.3% / 0.0 -100%
@@ 6 Results
Table 3: Evaluation of assumptions influence. CINIC-10 membership inference across five scenarios: Auxiliary-Knowledge Setting (all assumptions), removal of each assumption individually, and removal of all three simultaneously. Each entry shows TPR (%) at 0.01% / 0.0% FPR. The last column reports relative % of performance drop with respect to the Full Auxiliary-Knowledge Setting.
6.1 Membership Inference Results
We report membership inference attack results across three datasets, demonstrating the superiority of our proposed attack in the realistic scenario described above (without any assumptions on model training, data distribution, or member ratio). As shown in Tab. 1, reference-model attacks such as LiRA and RMIA struggle severely in this regime: at 0.0% False Positive Rate (FPR), LiRA achieves only 0.17% True Positive Rate (TPR) on CIFAR-10, and RMIA 0.01%. Importantly, these are the best-performing attacks, even compared to white-box methods, when the knowledge assumptions are provided (see Tab. T7 in the appendix). Their effectiveness relies on training large sets of reference models under matching conditions, a requirement that breaks down when the attacker cannot replicate the target’s training process. In contrast, ImpMIA avoids training reference models entirely, and maintains strong TPR at low FPR across all datasets. On CIFAR-10, our attack achieves 1.41% TPR at 0.0% FPR, and 2.76% at 0.01% FPR, with similar or larger gains on the other datasets. Our approach achieves substantially stronger performance at strict low-FPR operating points, which are the most relevant in practical membership inference scenarios, and comparable AUC results. ROC curves are shown for CIFAR-10 in Fig. 3 and for CIFAR-100 and CINIC-10 in Fig. S2. Importantly, our attack is computationally much cheaper than reference-based methods, as it does not require training reference models and is about 4× faster than reference-model attacks (see Sec. B.8). Finally, Tab. 2 shows that ImpMIA outperforms prior attacks on VGG16 and ResNet50 in the low-FPR regime.